) Find the value of integral Z C F · dr, where C is a semicircle parameterized by vecsr(t) = cost,sin t, 0 ≤ t ≤ π, and F =< −y, x > . 


1
Expert's answer
2021-11-03T11:32:24-0400

Given, F=<-y,x>=-y i+x j

and r=<cost, sint>=cost i+ sint j.

By using the formula,

"\int_C F.dr=\int_a^b F(r(t)).r'(t) dt\\
=\int_0^\pi F(cost, sint) .(-sint \space i+ cost \space
j) dt\\
=\int_0^\pi (-sint \space
 i+cost \space
j) .(-sint \space
 i+ cost \space
j) dt\\
=\int_0^\pi (sin^2t+cos^2t)dt\\
=\int_0^\pi dt\\
=[t]_0^\pi\\
=\pi"